### v2

# Personalized Generative Low-light Image Denoising and Enhancement

Abstract
Modern cameras produce remarkably high-quality images, yet their performance in low-light conditions remains suboptimal due to fundamental limitations in photon shot noise and sensor read noise. While generative image restoration methods have shown promising results compared to traditional approaches, they often suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low. Leveraging the availability of personalized photo galleries on users’ smartphones, we introduce Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that builds a customized diffusion model for individual users. Our key innovation lies in the development of an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery. This ID-consistent physical buffer serves as a robust prior that can be seamlessly integrated into the diffusion model to restore degraded images without the need for fine-tuning. Over a wide range of low-light testing scenarios, we show that DiffPGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
Figure 1: Using gallery from a phone user, we can restore low-light noisy facial images. Our method produces finer details and better identity consistency compared to other state-of-the-art restoration approaches like MIRNet [61, 62], GFP-GAN [52], DiffLL [22] and CodeFormer [67] on both simulated (top row) and real-captured (bottom row) low-light noisy inputs.
22footnotetext: Project page: https://genai-restore.github.io/DiffPGD/.
1 Introduction
With the astonishing development of smartphones, it is safe to say that smartphone cameras today have surpassed digital single-lens reflex (DSLR) cameras in both popularity and functionality. However, the small form factor of smartphone cameras puts a tight constraint on the aperture size, hence limiting the amount of light a CMOS pixel can detect. In a low-light imaging environment, this creates a fundamental limitation to how short the minimum exposure needs to be and how much signal-to-noise ratio (SNR) the sensor can support. While some of the noise seen at low light can be mitigated using better CMOS technology (e.g., correlated double sampling to reduce the read noise [41] and deep-well pump gate to reduce the stray capacitance between the floating diffusion and the transfer gate [39, 40]), the Poisson / Bose-Einstein statistics due to random photon arrivals is a problem created by mother nature that cannot be solved by even the ideal sensors. Therefore, image denoising and enhancement by exploiting the internal structures of images become a necessary task for all smartphone cameras.
From a signal-processing perspective, degradation caused by low light can be roughly approximated via
$\mathbf{y}=\text{ADC}\big{(}\text{Poisson}(\alpha\mathbf{x}+\eta)\big{)}+\epsilon,$ (1)
where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the sensor gain, $\eta$ is the dark current, and $\epsilon$ is the read noise. The inverse problem associated with low-light denoising is to recover $\mathbf{x}$ from $\mathbf{y}$. Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [36], total variation [44, 4], wavelets [9] to non-local means [3], BM3D [10], and recently convolution neural networks [25, 64, 50, 6, 2, 18, 57], transformers [35, 38, 63], and diffusion models [26, 60, 68]. A common theme across existing methods is to find and utilize a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images. While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images. Particularly in the context of human face recovery, the restored results often lack real identity and exhibit artifacts. This is attributed to the ill-posed nature of the problem and the lack of appropriate constraints.
Why Gallery Photos? Smartphone cameras today typically store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions. While these images have many variations, they are all about the same person(s). Therefore, if the imaging goal is to take a photo of this user, the gallery on the phone would be the best source to build a prior $p(\mathbf{x})$. This concept is illustrated in Fig. 2. In the context of diffusion-based image restoration, the initial solution space can be large since many candidate solutions are consistent with the noisy observation. The gallery provides a strong constraint to the search problem. This allows us to search for better quality images and preserve the person’s identity.
Figure 2: The restoration of inputs degraded by noise and low-light conditions is highly ill-posed. By incorporating additional high-quality gallery photos of the same identity, we significantly reduce the solution space, thereby achieving improved identity consistency in the restored images.
There are, however, two technical questions:
  • What class of images we should focus on: In this paper, we focus on human faces because of their relevance to phone users. We assume that the gallery photos have been selected and are reasonably informative. Product-level engineering such as pre-processing and selection of the gallery photos is important but beyond the scope of this paper.
  • How to incorporate the class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration? We do not want to train a restoration model from the scratch because the gallery can be small. We want to avoid fine-tuning as much as possible during inference. Moreover, the person’s identity needs to be preserved.
Physical Buffers to the Rescue. Given the gallery photos, what kind of prior information would be useful for restoration? Advancements in computer vision have made it possible to extract facial physical buffers—including albedo and normal maps—from a person’s gallery of images[43, 65, 47]. These physical buffers capture crucial identity-defining properties such as surface geometry and skin color, effectively encoding an individual’s unique identity. At the same time, it also eliminates the external influence of environmental lighting, pose, and other identity-independent variables. By incorporating this robust prior that disentangles the external influences, our approach can effectively extracts identity information from the gallery.
Contributions. Our contributions are threefold:
  1. 1.
    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to denoise low-light facial images while presering the identity.
  2. 2.
    Physical Identity Buffers: We propose the use of physical identity buffers from existing face galleries to enrich the restoration process, enabling accurate reconstruction of human faces under severe degradation.
  3. 3.
    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical and scalable to multiple users.
Extensive experiments demonstrate that our approach outperforms existing techniques in both visual quality and identity preservation, providing a significant advancement in low-light facial image restoration.
2 Related Works
Low-light Image Enhancement and Denoising. Low-light images often suffer from significant noise due to low signal-to-noise ratios (SNR), in addition to reduced brightness and contrast. Traditional low-light denoising techniques often focus on exploiting image structures, ranging from early methods like Wiener filtering [36], total variation [44, 4], and wavelets [9] to non-local means [3] and BM3D [10]. More recently, deep learning approaches, including convolutional neural networks [25, 64, 50, 6, 2, 18, 57, 66], transformers [35, 38, 63], and diffusion-based models [26, 60, 68, 42], have emerged as promising solutions for low-light denoising. Low-light image enhancement (LLIE) aims to improve the brightness and contrast of images captured in poorly lit conditions. Traditional approaches [16, 31, 29] often leverage heuristic or physical models to enhance image quality, while learning-based methods [28, 56, 24, 5, 17, 37, 53, 59, 22] have demonstrated greater effectiveness when supported by large-scale datasets. Recent advancements [45, 38, 42, 58] in this field have increasingly focused on jointly addressing denoising and enhancement tasks, aiming to improve image quality while effectively suppressing noise to maintain natural appearance. However, to the best of our knowledge, no specific approach currently exists for joint enhancement and denoising of facial images in low-light conditions.
Figure 3: Path to the idea. Using facial physical buffers and gallery images can help preserve human face identity information.
Deep Face Restoration. CNN/Transformer-based face restoration methods [54, 14, 23], such as RestoreFormer [54], typically rely on paired low- and high-quality image datasets to learn resolution enhancement and noise reduction. While effective, these methods often struggle under severe degradation and lack the capability to synthesize realistic details in complex scenarios.
In recent years, generative priors have been increasingly utilized in face restoration to better address real-world degradations. GAN-based methods, such as GFP-GAN [52], leverage StyleGAN2 [49] to enhance facial details, while diffusion-based models [55, 7, 13, 48, 51] have further expanded restoration capabilities. For instance, BFRDiffusion [7] uses Stable Diffusion to enhance low-quality images by adding high-fidelity details, and Ding et al. [13] employ a pretrained diffusion model that denoises degraded inputs while preserving identity by fine-tuning on a person’s gallery photos. MGBFR [48] combines text prompts and reference images with a dual-control adapter to retain accurate facial attributes and identity. Although these generative methods are effective, they often rely on multi-stage training or finetunning, which can limit the efficient use in some real-world applications.
Reference Prior for Face Restoration. In face restoration, reference images [48, 13, 32, 33] provide essential priors for recovering detailed and high-frequency facial features. These priors guide the model to generate realistic restorations. Reference-based methods can be divided into two types: single reference and gallery-based images. The single reference approach [48] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference. In contrast, the gallery-based approach [13] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration. Despite the effectiveness [13] using gallery photos of users, it still requires multi-stage training and fine-tuning on each user during inference.
3 Method
Figure 4: For images affected by mild low-light noise, phycial buffers can be accurately extracted. However, when encountering the images degraded by severe low light and noise, precise physical buffers are difficult to obtained.
3.1 Restoration by Diffusion
A diffusion model approximates the clean image distribution $p_{\theta}(\mathbf{x}_{0})$ by learning a model $\theta$ that reverses the noise-adding process. In Denoising Diffusion Probabilistic Models (DDPMs) [21], Gaussian noise is progressively added to a clean image $\mathbf{x}_{0}$:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\text{where}\quad\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (2)
The reverse process then removes the noise to reconstruct the clean image. After training, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, the model can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t})$.
In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $\mathbf{x}_{0}$ from a low-quality input $\mathbf{y}$. Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $\mathbf{y}$. During training, the model uses $\mathbf{y}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, it can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})$ using the trained model.
Figure 5: The overall architecture of the proposed method. Our core idea is to use ID-consistent physical buffers, extracted from gallery photos, to constrain the generative space in the diffusion model restoration process. For a high-quality gallery, we use LAP [65] to extract the albedo and normal information for each photo and apply adaptive aggregation to fuse the entire gallery. The extracted albedo represents base skin color and facial appearance, while normal captures facial geometry. In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
3.2 Why Combining Gallery and Physical Buffer?
We formalize our core approach of extracting identity-consistent physical buffers from gallery photos to guide the diffusion model for restoration. This method is based on the observation that using facial physical buffers or gallery images can help preserve identity information during restoration tasks involving diffusion models. To evaluate the capability of physical buffers and gallery images, we conducted four simple experiments under a blind scheme (excluding $y_{0}$ during training), as illustrated in Fig. 3:
  • Exp 1: Model trained solely on generic images (purely generative).
  • Exp 2: Model trained on generic images, conditioned on the facial physical buffer extracted from the input image.
  • Exp 3: Model trained on generic images, then fine-tuned with the user’s gallery photos.
  • Exp 4: Model trained on generic images, fine-tuned with the user’s gallery photos, and conditioned on the facial physical buffer.
We tested these models on a mildly noisy input. Exp 1 produces random faces, as expected from a purely generative model. In contrast, Exp 2 uses physical buffers obtained via DECA [15], which can extract coarse face buffers from a single 2D photo (see Fig. 4), yielding faces with user-like features (e.g., nose shape, mouth and skin tone). Hence, physical buffers can convey key identity cues. Exp 3 outputs distinct faces of the same user, compared to Exp1, showing that fine-tuning with gallery photos effectively transfers identity information. Finally, Exp 4 yields the best results, showcasing that the collaborative use of gallery photos and facial physical prior can achieve the optimal outcomes.
While Fig. 3 shows us a design direction, we need to overcome:
  1. (i)
    Fine-tuning: Extracting identity via fine-tuning requires significant time and must be repeated for each user. Additionally, the small size of a gallery compared with a large pre-trained model can lead to overfitting.
  2. (ii)
    Unreliable buffers from degraded inputs: As studied in the face reconstruction literature, low-quality inputs consistently present additional challenges [15]. Severe degradation undermines the extraction of physical buffers (Fig. 4), resulting in poor reconstruction. Moreover, features captured from a single degraded photo are limited, not to mention from a degraded one.
In conclusion, to avoid repeated fine-tuning on galleries and reduce the risk of erroneous buffers from noisy inputs, it is more desirable to jointly employing gallery photos and facial physical priors, as we will discuss below.
3.3 Extracting Physical Buffer from Photo Gallery
Based on the preceding analysis, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training. This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing more acurate facial identity features of the target person. Specifically, We will extract the albedo and normal maps of the target person, which are two key attributes used to describe the physical characteristics of a face.
  • Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects. It reveals a person’s facial appearance, e.g. eyebrows, lip coloration, and other distinct texture elements can reflect individual identity traits.
  • Normal encodes the surface orientation at each pixel, which provides hints about the underlying shape characteristics.
By conditioning on these identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction. Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
However, extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA. This complexity arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses. To ensure these buffers comprehensively represent the person’s identity and facial properties, they must be aligned to maintain identity consistency. As shown in the physical buffer extractor module in Fig. 5(a), we adopt the aggregation net design from the LAP [65] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
3.4 Model Framework
We name our final proposed model DiffPGD. As illustrated in Fig. 5, the model leverages ID physical buffers (albedo $a^{i}$ and normal $n^{i}$) extracted from person i’s gallery photos as conditions to guide our diffusion model during training directs the generative process toward a personalized space.
The aggregation network contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting the consistent representation. To model albedo and normal, two separate aggregation networks, denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed. Given a photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $\mathbf{\alpha}_{k}^{a}$ and $\mathbf{\alpha}_{k}^{n}$, respectively. To derive a global representation of the identity for person i based on $\{\mathbf{\alpha}_{k}^{a},\mathbf{\alpha}_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{\mathbf{w}_{k}^{a},\mathbf{w}_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$. The combined global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ for texture and depth are calculated as:
$\mathbf{\alpha}_{c}^{a}=\sum_{k=1}^{N}\mathbf{w}_{k}^{a}\mathbf{\alpha}_{k}^{a},\text{ and}\quad\mathbf{\alpha}_{c}^{n}=\sum_{k=1}^{N}\mathbf{w}_{k}^{n}\mathbf{\alpha}_{k}^{n}.$ (3)
Once calculated, the global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ are passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$.
During inference, the process begins with a Gaussian noise image $\mathbf{x}_{T}^{i}$, which is iteratively transformed into the clean, enhanced image $\mathbf{x}_{0}^{i}$ by sampling from $p(\mathbf{x}_{0}^{i}|\mathbf{x}_{T}^{i},\mathbf{y}^{i},\mathbf{a}^{i},\mathbf{n}^{i})$ using the trained model. With the low-light noisy image $\mathbf{y}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
Our experiments demonstrate that identity information is effectively preserved with the guidance of these ID-consistent buffers, enabling our model to achieve superior results in human face restoration.
4 Experiments
Figure 6: Visual comparison on real cases. The first column presents the input images along with their capture parameters. In the second column, we enhance the brightness to better reveal the noise. Please zoom in for the best visual experience.
4.1 Dataset
Training Data. Since no publicly available datasets exist for low-light noisy face images, we simulate such data using the CelebAMask-HQ dataset [30], the InverseISP algorithm [34], and the Poisson-Gaussian noise model Eq. 1. The pipeline involves converting RGB images to their corresponding RAW mosaiced format using InverseISP [34], followed by applying Eq. 1 to simulate low-light noise using realistic camera sensor parameters. Multiple noise levels are generated by varying the light intensity (Lux) to ensure the model is robust across diverse noise conditions. Refer to supplementary for more details on the simulation.
  Photons Per Pixel 5 10 13 26
Methods ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$
SNR-aware[59] 0.2156 188.3311 0.1302 8.4541 0.4240 139.8193 0.0786 9.5462 0.4999 124.1162 0.0626 10.1712 0.6937 99.9875 0.0383 12.4961
MIRNet[61, 62] 0.1229 294.7381 0.2414 8.6008 0.2860 214.7785 0.1534 9.8572 0.3830 187.2571 0.1286 10.4209 0.6212 136.1320 0.0793 12.8992
DiffLL[22] 0.2910 167.3905 0.1068 17.3506 0.4544 150.4559 0.1117 18.3360 0.5095 140.2738 0.0996 19.9547 0.6833 119.2693 0.0760 23.3180
GFPGAN[52] 0.1941 388.9293 0.4256 7.1990 0.2075 233.3173 0.1521 8.1283 0.2976 190.7264 0.1326 9.5483 0.5976 158.9475 0.0829 11.4855
CodeFormer[67] 0.2863 111.4544 0.04908 18.2406 0.4940 91.1244 0.0326 19.2692 0.5563 84.7955 0.0281 20.2306 0.7338 80.6582 0.0263 24.5512
Ours 0.3714 77.5743 0.0084 17.9334 0.5225 71.8645 0.0063 18.4999 0.5938 69.5567 0.0042 19.0910 0.7118 59.1975 0.0030 23.7226
Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset on 5PPP, 10PPP, 13PPP, and 26PPP. The lower the PPP, severe the degradation is.
Testing Data. For testing DiffPGD, we utilize both simulated and real-world captured low-light noisy data.
The simulation setup assumes a $12$-bit image sensor with parameters similar to that of Canon DSLR cameras: a Full Well Capacity of $36000\text{e}^{-}$, Quantum Efficiency of $0.42$, dark current of $7\text{e}^{-}$, and read noise of $6\text{e}^{-}$. These parameters ensure realistic noise characteristics for evaluating model performance.
We also captured 100 RAW-format face images as real-world test dataset using a Sony $\alpha 6400$ DSLR camera paired with a Sony 18-135mm lens at night. These images feature diverse low-light and dark backgrounds, various angles and poses, as well as different focal lengths, apertures, ISO values, and shutter speeds to reflect the rich characteristics of real-world low-light photography. We cropped these images into 256×256 face patches.
4.2 Implementation Details
Our denoising model is built on the ADM architecture [12], and the ID-consistent physical extractor is implemented using the aggregation network from LAP [65]. For training, we use a P2-weighted loss function [8], which measures the discrepancy between the predicted and ground-truth noise. This loss is defined as: $L=\lambda_{t}||\hat{\epsilon}_{t}-\epsilon_{t}||_{2}^{2}$ , where $\lambda_{t}$ serves as a hyperparameter to adjust the loss weight dynamically across different timesteps.
During the training stage, we train our model on the CelebAMask-HQ dataset [30], comprising 30,000 images distributed across 4,516 unique identities(IDs). We use 3318 IDs for training and 200 IDs for simulated testing. Within each ID, there are 3–20 images. The number of gallery images we used per person varies from 3 to 6, with most around 5. The Adam optimizer is applied with a learning rate of 1e-4. Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model.
4.3 Performance in Real-Captured Scenarios
Testing model performance in real-life low-light scenarios is crucial and challenging, as real-world conditions often involve uncontrollable factors compared to simulated ones. Since no existing data set addresses the restoration of real-life low-light noisy face photos, we captured photos of real persons under indoor and outdoor low-light scenarios , as illustrated in Section 4.1.
We compare our methods with several state-of-the-art works: UTV-Net[66], GFPGAN[52], SNR-aware[59], MIRNet[61, 62], DiffLL[22], Codeformer [67]. These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research. Particularly, GFPGAN and Codeformer are designed specifically for face image restoration, which is highly competitive with our method. We re-trained/fine-tuned all the baselines with the same training dataset as ours for fair comparison.
Quantitative and qualitative results are presented. To assess the quality of restored images, we employ the Fréchet Inception Distance (FID) [20] and the Kernel Inception Distance (KID) [1], as these metrics offer a robust evaluation aligned with human perception. Since ground-truth images are unavailable for real-captured cases, we compute the FID and KID score between the model’s output and the corresponding subjects’ gallery photos. As shown in Fig. 6, our method produces images with the highest quality and level of detail. The restored images exhibit more natural skin tones and well-preserved facial identity features, such as eye shape, eyebrows, and lips, while maintaining the highest consistency in facial identity.
Since our work focuses on face images, we also check on the identity preservation performance. Besides the image quality metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[11] to further evaluate the identity consistency. Again, since ground truth images are unavailable for real-captured cases, we calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo of the corresponding subjects. The results are also shown in Table 2. Again, our method has the highest identity score, and the visual results are consistent with the quantitative metric results. The results demonstrate the supreme ability of our proposed DiffPGD model on real low light denoising and enhancement task.
4.4 Performance in Simulated Scenarios
In simulated scenarios, We use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set. We simulated various low-light noisy levels by setting different photons per pixel (PPP), i.e., the average number of photons arriving at the sensor within the exposure time. The lower the PPP, severe the degradation is. A lower photon level increases the impact of read noise and dark current on the final output, thereby decreasing the quality of the reconstructed images.
Quantitative results are provided in Table 1. The image restored from our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistent best. (Refer to supplementary for more visual comparisons). From the table, we could see that CodeFormer surpasses us on PSNR. Fig. 7 illustrates a visual comparison between our method and CodeFormer under inputs ranging from 5 to 65 PPP. We observe that CodeFormer tends to generate smoother images while we could preserve better high-frenquency details. Furthermore, our method demonstrates superior identity preservation even at low resolutions, such as 5 PPP to 13 PPP, while exhibiting fewer artifacts. The advantage of our model in maintaining identity consistency is particularly evident in heavily degraded cases. The results highlight the strong performance of our proposed DiffPGD model in low-light denoising and enhancement tasks.
  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
SNR-aware[55] 252.2397 0.1264 0.1990
UTVNet[66] 341.8502 0.1617 0.2002
MIRNet[61, 62] 276.3853 0.1659 0.2210
DiffLL[22] 252.1553 0.1421 0.3196
GFPGAN[52] 252.7411 0.1249 0.1525
CodeFormer[67] 281.9744 0.1459 0.2965
Ours 251.9318 0.0882 0.3708
Table 2: Comparisons with state-of-the-art works on the real-captured test cases.
  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
w/o physical buffers 272.6148 0.1281 0.3169
w/ physical buffers 251.9318 0.0882 0.3708
Table 3: Ablation comparisons of using ID physical buffers on real-captured photo test.
Figure 7: Visual comparison on the simulated cases. Our method maintains good identity consistency even at extremely low PPP and exhibits fewer artifacts.
Figure 8: The relationship between the identity (ID) score and FID of the restored images and the number of gallery images used to generate the ID physical buffer.
5 Ablation Study
ID-Consistent Physical Buffers for Identity Preservation. We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-captured testing scenarios. In Table 3, we compare the quantitative results of training our model with and without ID physical buffers in real-captured test cases. The results demonstrate that using physical buffers largely increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
Effect of the Number of Gallery Images. The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the User’s gallery photos. As claimed in the paper, facial features extracted from a single image are inherently limited. In Fig. 8, we evaluate the effect of the number of gallery images used to generate the ID physical buffer. We analyze it with the real-captured set where 6 gallery images were collected from each subject. The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image. Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case without gallery photos.
6 Conclusion
In this work, we addressed the challenges of low-light facial image restoration by introducing Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that leverages user-specific photo galleries to improve denoising performance. By incorporating an identity-consistent physical buffer, we provided a strong prior that enhances the diffusion model’s ability to restore degraded images while avoiding the need for model fine-tuning. Our experiments, which included extensive evaluations on multiple challenging datasets, demonstrated that DiffPGD not only outperforms existing diffusion-based denoising methods but also delivers superior quantitative and qualitative results across various testing scenarios. This approach effectively mitigates false content generation in low-SNR conditions while bridging the gap in low-light imaging quality, ultimately paving the way for personalized and more robust image restoration techniques.
Supplementary Material
Figure 9: Simulation process for generating low-light noisy face images.
Appendix A Dataset
As mentioned in Section $4.1$ of the main paper, since no open-sourced low-light noisy datasets are available, we adopt a simulation approach to generate low-light noisy face images for training our model. Specifically, we utilize the CelebAMask-HQ dataset’s [30] face images as the ground truth RGB images.
As shown in Fig. 9, the first step in the simulation process involves passing the RGB image through a pre-trained Inverse ISP network [34] to generate a pseudo ground truth RAW mosaic image. For visualization purposes, we display the demosaiced version of this image. We utilize the pretrained weights provided by the authors [34], which were obtained by training the network on iPhone camera images, for unprocessing the RGB images into RAW format.
This pseudo ground truth RAW image is then passed through a Poisson-Gaussian Simulator to simulate the low-light noisy RAW mosaic image. It is noteworthy that during the training process, we vary the camera sensor parameters (in the Poisson-Gaussian simulator) across a wide range of values to enable the model to handle diverse camera sensor noise profiles. The simulation setup assumes a $12$-bit image sensor with parameters similar to those of Canon DSLR cameras. The parameter ranges are summarized in Table 5.
Table 5: Simulation parameters used in the Poisson-Gaussian Simulator.
  Parameter Range
Full Well Capacity $[19000\text{e}^{-},64000\text{e}^{-}]$
Quantum Efficiency $[0.32,0.54]$
Dark Current $[2.2\text{e}^{-},11.7\text{e}^{-}]$
Read Noise $[2.2\text{e}^{-},10.8\text{e}^{-}]$
Photons-per-Pixel (PPP) $[13\,\text{PPP},65\,\text{PPP}]$
These parameters are carefully chosen to ensure realistic noise characteristics during training.
Appendix B Additional Visual Results
In Fig. 12, Fig. 14, and Fig.15, we present additional visual comparisons between our model and baseline methods on real-captured and simulated cases. Our results consistently demonstrate superior performance, particularly in cases of severe degradation.
Figure 10: A network architecture diagram for identity feature restoration using a face encoder.
  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
w/ Encoder 12.62 0.2574 0.6315 0.6906
w/ physical buffers 13.85 0.3753 0.6042 0.8039
Table 6: Comparison of ID physical buffers and encoder-based methods on the real-captured photo test set. For a fair comparison, we use two same gallery images for each individual for both the encoder-based and physical buffer-based approaches.
Appendix C More Implementation Details
During training and testing, the image size is 256 $\times$ 256. The architecture of the diffusion model is adapted from [12]. We modify the architecture so that the model can take the concatenation of physical buffers and degraded images as conditions. The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000, and the two hyperparameters P2_gamma and p2_k are both set to 1.0. For additional architectural details, refer to [12]. During inference, we adopt the DDIM method [46] with 100 denoising steps. The model is trained using the Adam optimizer [27] with a learning rate of $10^{-4}$. The weight decay is set to 0. All experiments are performed on a single NVIDIA A100 GPU.
Appendix D Encoder vs. Physical Buffer?
Many previous studies have used encoder-based methods to extract latent facial feature information from reference photos[51]. To evaluate whether the encoder can effectively extract enough facial features to guide the diffusion model in solving the low-light enhancement and denoising task, we conducted an experiment. As illustrated in Fig. 10, we utilized an encoder (ResNet-18 [19]) to extract latent codes from the target individual’s gallery photos. Following the method employed by ADM [12] for time embedding, we scaled and shifted the features in each layer of the diffusion model using the latent codes. This process integrates the identity information extracted by the encoder into the diffusion model.
We compared the restored images generated with the encoder-based method with the proposed physical buffer-guided method on our real-captured photo set. This is a special semi-real set, in which we take photos of some printed paper photos at night. We prepare this semi-real set because The ground-truth will be available to be compared with. The quantitative and visual comparisons are presented in Table 6 and Fig. 11, respectively. The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the encoder-based approach. One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery. In contrast, the encoder-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle. Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
Appendix E Gallery Photos
Fig. 13 presents the the gallery photos for the test subjects included in the Fig. 6 of the main paper.
Figure 11: Visual comparison on different identity representation strategies (physical buffer and face encoder).
Figure 12: Additional real-captured visual comparison.
Figure 13: Gallery photos.
Figure 14: Additional visual comparison.
Figure 15: Additional visual comparison.